The most recent financial statements for Bello Co. are shown here: Income Statement Sales Costs Taxable Balance Sheet Current $12.000 Debt $16,600 14,000 Fixed assets 32,850 Equity 28,250 assets $ 6,400 Total $44,850 Total $44,850 income Taxes (2296) 1,408 Net income $ 4,992 Assets and costs are proportional to sales. Debt and equity are not. The company maintains a constant 35 percent dividend payout ratio. What is the internal growth rate? (Do not round intermediate calculations and enter your answer as a percent rounded 2 decimal places, e.g., 32.16.) Internal growth rate
Answer #1

Internal Growth Rate = (ROA x b) / [1-(ROA x b)]

ROA = Net Income / Total Assets

=4992/44850=11.13%

the plowback ratio, b, which is one minus the payout ratio. b = 1 - .5 = 65%

Internal Growth Rate = (11.13% x 65%) / [1-(11.13% x 65%)]=7.80%